Calculate the equivalent wt od N2 when converted to N2H4
 

Grade:11

1 Answers

Arun
5 years ago

Molecular Weight of Hydrazine (N2) is 28

Equivalent weight = 28/2 = 14 (due to presence of two lone pairs its acidity is 2.)